Write a function using this form. y=ab^x and these points (2,56)(0,3.5)
Answer by Nate(3500)

y = a*b^x
3.5 = a*b^0
3.5 = a
~~~~~>
y = 3.5*b^x
56 = 3.5*b^2
16 = b^2
4 = b
~~~~~>
y = 3.5*(4^x)